OpenStudy (anonymous):

Which equation best represents the situation? A teacher prepared math problems for students to solve in groups. She split her class into 4 groups. Each group got an equal number of problems. After distributing them, there were 2 left over. Group A completed their problems plus 2 extras (30 problems). How many problems did the teacher originally prepare?

OpenStudy (anonymous):

A. 4x+2=30 B.x/4+2=30 C.x/2-4=30 D.2x-4=30